OpenStudy (anonymous):

WILL MEDAL Choose the answer. What is the solution to the equation? A.z = –44 B.z = –20 C.z = 20 D.z = 44

OpenStudy (jessereyna):

Can you post the equation?

OpenStudy (anonymous):

\[3-\frac{ z }{ 4 }= 8\]

OpenStudy (landon34):

I can solve it.

OpenStudy (jessereyna):

Okay you need to first subtract three from each side

OpenStudy (jessereyna):

Then multiply 4

OpenStudy (anonymous):

i just need the answer

OpenStudy (landon34):

z in (-oo:+oo) 3-(z/4) = 8 // - 8 3-(z/4)-8 = 0 -1/4*z-5 = 0 // + 5 -1/4*z = 5 // : -1/4 z = 5/(-1/4) z = -20 z = -20

OpenStudy (anonymous):

thanks

OpenStudy (jessereyna):

B

OpenStudy (landon34):

your welcome.